OpenStudy (anonymous):

Write the expression in standard form. 5/(3-15i)

OpenStudy (jdoe0001):

multiply the numerator and denominator by the denominator's conjugate $$ \cfrac{5}{3-5i} \times \cfrac{3+5i}{3+5i} $$

OpenStudy (jdoe0001):

keep in mind the difference of squares, \((a-b)(a+b) = (a^2-b^2)\)

OpenStudy (anonymous):

is the answer \[\frac{ 5 }{ 78} + \frac{ 25 }{ 78 }i\]

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

so is this the right answer?

OpenStudy (jdoe0001):

well $$ \cfrac{5}{3-5i} \times \cfrac{3+5i}{3+5i}\\ \implies \frac{5(3+5i)}{3^2-(5i)^2} \implies \frac{15+25i}{9-25(i^2)} $$

OpenStudy (jdoe0001):

wait, you have 15i, lemme rewrite that using 15i, NOT 5i

OpenStudy (jdoe0001):

$$ \large { \cfrac{5}{3-15i} \times \cfrac{3+15i}{3+15i}\\ \implies \frac{5(3+15i)}{3^2-(15i)^2} \implies \frac{15+75i}{9-225(i^2)} } $$

OpenStudy (anonymous):

yes after this is where i got lost

OpenStudy (jdoe0001):

now, \(i^2 = (\sqrt{-1})^2 \implies -1\)

OpenStudy (anonymous):

i^2 is -1 so it would be \[\frac{ 15+75i }{ 9+225 }\]

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

\[\frac{ 15+75i }{ ?234}\]

OpenStudy (jdoe0001):

yes, can further simplify it by getting common factor of "3" above and below

OpenStudy (anonymous):

ok thank you

OpenStudy (jdoe0001):

$$ \large{ \cfrac{ \cancel{3}(5+25i) }{ \cancel{3}(78) } } $$

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

or 5/78 + (25/78)i